The sum of t
hree numbers is five. The first number minus the second number plus the third number is one. The first number minus the third number is three more than the second number. Find the numbers

Answer by stanbon(48564)

I'll set up the equations and let you solve the system.
Let the numbers be a, b, c.
Then:
a+b+c=5
a-b+c=1
a-c=b+3
Cheers,
stan H.